OpenStudy (anonymous):

hi, how many principles of differentiation is there? can anyone please state them (with formulas) for me?

OpenStudy (anonymous):

I can help what subject is this?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

okay first tell me what you know

OpenStudy (anonymous):

nvm there are 5 principles

OpenStudy (anonymous):

wow...ok

OpenStudy (anonymous):

i know differentiation and integration in first principle

OpenStudy (anonymous):

i never knew of the other principles

OpenStudy (anonymous):

i know how to find the various derivatives and integrals...thats all.

OpenStudy (anonymous):

I didn't either lol but its quite simple www.pkwy.k12.mo.us/candd/.../gifted/PrinciplesofDifferentiation.htm this can help you

OpenStudy (anonymous):

okay...I'd try it out and get back to you

OpenStudy (anonymous):

Thanks for the help

OpenStudy (unklerhaukus):

\[.\newcommand \de [2] {\displaystyle\frac{\mathrm d #1}{\mathrm d#2}} % first order derivative\] CONSTANT RULE The derivative of a constant equals zero. \[\boxed{\de{}x(c)=0\qquad c\in \mathbb R}\]\[c' = 0\] POWER RULE The derivative of a power of \(x\), equals the power multiplied by \(x\) to the power reduced by one. \[\boxed{\de{}x(x^n) = nx^{n-1}}\] SUM RULE The derivative of a sum, equals the sum of the derivatives. \[\boxed{\de{}x(f + g) = \de fx + \de gx}\]\[(f + g)' = f' + g'\] PRODUCT RULE The derivative of a product equals, the the derivative of the first term multiplied by the second term, plus the first term multiplied by the derivative of the second term. \[\boxed{\de {}x(f\cdot g) = \de fx\cdot g + f\cdot\de gx}\]\[(f\cdot g)' = f'\cdot g + f\cdot g'\] QUOTIENT RULE The derivative of a quotient, equals the derivative of the numerator multiplied by denominator, minus the numerator multiplied by the derivative of the denominator, all divided by the square of the denominator. \[\boxed{\de{}x\Big(\frac fg\Big) = \frac{\de fx\cdot g-f\cdot\de gx}{g^2}}\] CHAIN RULE The derivative of a composition, equals the composition of the derivative of the first term with the second term, multiplied by the derivative of the second term. \[\boxed{\de {}x(f\circ g) = \de fg\cdot\de gx}\]\[\big(f\circ g\big)' = (f'\circ g)\cdot g'\]